OpenStudy (anonymous):

the equation 2w + 5j =60 is used to determine the number of water bottles w and the number of juice bottles j that can be brought for $60 .if you purchase 4 bottles of juice how many bottles of water can u buy a.10 b.15 c.20 d.25 can you explain an show how to do this problem

OpenStudy (anonymous):

Sorry, 20

OpenStudy (anonymous):

ok. From the equation, it can be inferred that each water bottle costs 2$ and each juice bottle costs 5$ . When you purchase 4 juice bottles, you just plug in 4 in place of j in the equation. Which gives you: 2w + 5*4 = 60 2w = 60 - 20 = 40 Therefore, w = 20 Therefore, number of water bottles that can be purchased = 20